# Matteo Escudé

### v5

@@ A.2.2 Existence proof
If $s_{-}=s_{+}=\overline{p}$ then it must be that $\operatorname*{cav}w_{\beta}(\overline{p})=w_{\beta}(\overline{p})$, by definition of $s_{-}$ and $s_{+}$. Suppose otherwise that $\overline{p}$ is not lowest-consistent. Then there exists some $\tilde{s}<\overline{p}$ such that $w_{\beta}(\tilde{s})\geq w_{\beta}(\overline{p})$, since any message available at $\overline{p}$ (including optimal ones) is also available at some lower type. Recall also that, since the sender can prove news better than the prior, there exists some message $m_{h}$ such that $w_{\beta}(\min M^{-1}(m_{h}))>w_{\beta}(\overline{p})$ with $\min M^{-1}(m_{h})>\overline{p}$. Summarizing, we have that $\tilde{s}<\overline{p}<\min M^{-1}(m_{h})$ with $w_{\beta}(\tilde{s})\geq w_{\beta}(\overline{p})$ and $w_{\beta}(\min M^{-1}(m_{h}))>w_{\beta}(\overline{p})$. Since $\operatorname*{cav}w_{\beta}$ is concave and majorizes $w_{\beta}$, it must be that $(\operatorname*{cav}w_{\beta})(\overline{p})>w_{\beta}(\overline{p})$, contradicting that $(\operatorname*{cav}w_{\beta})(\overline{p})=w_{\beta}(\overline{p})$.
Claim 2.
If $s_{-}<\overline{p}<s_{+}$, then $s_{-}$ and $s_{+}$ are lowest-consistent.
Proof.
We make use of the following intermediate claim, proven after the main argument is complete.
@@
Suppose first that $s_{-}$ is not lowest-consistent. Then there exists a $\tilde{s}<s_{-}$ such that $w_{\beta}(\tilde{s})\geq w_{\beta}(s_{-})$. Since $\operatorname*{cav}w_{\beta}$ majorizes $w_{\beta}$ we have that $(\operatorname*{cav}w_{\beta})(\tilde{s})\geq w_{\beta}(\tilde{s})$ and, by definition of $s_{-}$, $(\operatorname*{cav}w_{\beta})(s_{-})=w_{\beta}(s_{-})$. It follows that $(\operatorname*{cav}w_{\beta})(\tilde{s})\geq(\operatorname*{cav}w_{\beta})(s_{-})$. Observe next that Claim 3 and the definition of $s_{-}$ and $s_{+}$ imply that $(\operatorname*{cav}w_{\beta})(s_{+})>(\operatorname*{cav}w_{\beta})(s_{-})$. Summarizing, we have that $\tilde{s}<s_{-}<s_{+}$ with $(\operatorname*{cav}w_{\beta})(\tilde{s})\geq(\operatorname*{cav}w_{\beta})(s_{-})$ and $(\operatorname*{cav}w_{\beta})(s_{+})>(\operatorname*{cav}w_{\beta})(s_{-})$, contradicting that $\operatorname*{cav}w_{\beta}$ is concave.
Suppose next that $s_{+}$ is not lowest-consistent. Then there exists a $\tilde{s}<s_{+}$ such that $w_{\beta}(\tilde{s})\geq w_{\beta}(s_{+})$. Observe next that the definition of $s_{-}$ and $s_{+}$ and Claim 3 imply that $(\operatorname*{cav}w_{\beta})(s_{+})=w_{\beta}(s_{+})>w_{\beta}(s_{-})=(\operatorname*{cav}w_{\beta})(s_{-})$. Since $w_{\beta}(\tilde{s})\geq w_{\beta}(s_{+})$, we have that
@@
There are three possibilities. If $\tilde{s}<s_{-}$, then $\operatorname*{cav}w_{\beta}$ is not concave, a contradiction. If $\tilde{s}=s_{-}$ then $(\operatorname*{cav}w_{\beta})(s_{-})>(\operatorname*{cav}w_{\beta})(s_{-})$, also a contradiction. If $\tilde{s}\in(s_{-},s_{+})$ then $\operatorname*{cav}w_{\beta}$ is not affine over $[s_{-},s_{+}]$, contradicting Lemma 2. It follows that both $s_{-}$ and $s_{+}$ are lowest-consistent, completing the proof of the claim.
@@
Suppose otherwise that $w_{\beta}(s_{-})\geq w_{\beta}(s_{+})$. Since the sender can prove news better than the prior, there exists a message $m_{h}$ such that $v(\min M^{-1}(m_{h}))>v(\overline{p})$. Note that $\min M^{-1}(m_{h})>\overline{p}$, since $v$ is non-decreasing. It follows from the definition of $w_{\beta}$ that $w_{\beta}(\min M^{-1}(m_{h}))>v(\overline{p})$.
Observe next that $w_{\beta}(s_{-})\leq v(\overline{p})$, since for any $m\in M(s_{-})$, $\beta(m)\leq s_{-}$ and therefore $v(\beta(m))\leq v(s_{-})\leq v(\overline{p})$. This, in turn, implies that $w_{\beta}(\min M^{-1}(m_{h}))>w_{\beta}(s_{-})$. Hence, since we supposed that $w_{\beta}(s_{-})\geq w_{\beta}(s_{+})$, we have that $w_{\beta}(\min M^{-1}(m_{h}))>w_{\beta}(s_{+})$ also holds. Summarizing, we have established that $w_{\beta}(\min M^{-1}(m_{h}))>w_{\beta}(s_{-})\geq w_{\beta}(s_{+})$.
Recall that, by definition, $(\operatorname*{cav}w_{\beta})(s_{-})=w_{\beta}(s_{-})$ and $(\operatorname*{cav}w_{\beta})(s_{+})=w_{\beta}(s_{+})$. There are three possibilities. If $\min M^{-1}(m_{h})<s_{+}$, $\operatorname*{cav}w_{\beta}$ is not affine over $(s_{-},s_{+})$, contradicting Lemma 2. If $\min M^{-1}(m_{h})=s_{+}$ then $w_{\beta}(s_{+})>w_{\beta}(s_{+})$, a contradiction. If $\min M^{-1}(m_{h})>s_{+}$ then, since
$(\operatorname*{cav}w_{\beta})(\min M^{-1}(m_{h}))>(\operatorname*{cav}w_{\beta})(s_{-})\geq(\operatorname*{cav}w_{\beta})(s_{+}),$
$\operatorname*{cav}w_{\beta}$ is not concave, a contradiction. It follows that $w_{\beta}(s_{-})<w_{\beta}(s_{+})$.
This completes the proof of Lemma 4.
@@ Appendix B Proof of Theorem 1
Suppose otherwise: there exists an equilibrium $(\pi,\mu,\beta)$ with $s^{\prime\prime}\in\operatorname*{supp}\pi_{S}$ and a message $m^{\prime\prime}\in M(s^{\prime\prime})$ such that $\mu(m^{\prime\prime}|\pi,s^{\prime\prime})>0$ and such that for some $s^{\prime}<s^{\prime\prime}$, $m^{\prime\prime}\in M(s^{\prime})$.
Consider first the case in which the sender acquires no information in equilibrium, so that $\pi_{S}(\overline{p})=1$ and $s^{\prime\prime}=\overline{p}$. The on-path receiver belief must be correct: $\beta(m)=\overline{p}$ for all $m\in\operatorname*{supp}\mu(\cdot|\pi,\overline{p})$. The sender’s equilibrium payoff is therefore $v(\overline{p})$. Because the sender can prove news better than the prior, there exists some $m_{h}$ such that $v(\min M^{-1}(m_{h}))>v(\overline{p})$ and, since $v$ is non-decreasing, we also have that $s_{h}\equiv\min M^{-1}(m_{h})>\overline{p}$. We will show that the sender can profitably deviate by choosing signal $\pi^{\prime}$ with $\operatorname*{supp}\pi_{S}^{\prime}=\{s^{\prime},s_{h}\}$. Following realization $s^{\prime}$ the sender obtains a payoff of at least $v(\overline{p})$ (since $m^{\prime\prime}\in M(s^{\prime})$, by construction) and following realization $s_{h}$ the sender obtains a payoff strictly above $v(\overline{p})$ (since $m_{h}\in M(s_{h})$, $\beta(m_{h})\geq s_{h}$ and therefore $v(\beta(m_{h}))\geq v(s_{h})>v(\overline{p})$). Hence $\pi^{\prime}$ is a strictly profitable deviation, contradicting that $(\pi,\mu,\beta)$ is an equilibrium.
Consider next the case in which the sender does acquire information in equilibrium, so that $\operatorname*{supp}\pi_{S}\neq\{\overline{p}\}$. Define the sender’s equilibrium interim value from signal realization $s\in[0,1]$ as $w_{\beta}(s)\equiv\max_{m\in M(s)}v(\beta(m))$. The argument relies on the following claim, proven in Appendix B.1.
@@
Let $\underline{s}\equiv\min\operatorname*{supp}\pi_{S}$ and $\overline{s}\equiv\max\operatorname*{supp}\pi_{S}$. Since the sender acquires information, $\underline{s}<\overline{p}<\overline{s}$. Suppose first that $s^{\prime\prime}\leq\overline{p}$. Since $w_{\beta}$ is affine (by Claim 4) and $\sum_{s\in\operatorname*{supp}\pi_{S}}\pi_{S}(s)s=\overline{p}$, the equilibrium payoff can be written as
@@
Consider now a deviating signal supported on $\{s^{\prime},\overline{s}\}$ only. By choosing this signal, the sender can obtain at least a payoff of
@@
To show that (7) is strictly larger than (6) it suffices to note that: $w_{\beta}(s^{\prime})\geq w_{\beta}(s^{\prime\prime})$, since $m^{\prime\prime}\in M(s^{\prime})$ by assumption; $s^{\prime}<s^{\prime\prime}$, by assumption; and $w_{\beta}(\overline{s})>w_{\beta}(s^{\prime\prime})$, since $w_{\beta}$ is strictly increasing on $\operatorname*{supp}\pi_{S}$ (by Claim 4). This contradicts that $\pi$ is an equilibrium signal.
@@
Suppose next that $s^{\prime\prime}>\overline{p}$. There are three cases and, in each, a deviation analogous to the one in the previous paragraph can be constructed: (i) if $s^{\prime}>\overline{p}$, a signal supported on $\{\underline{s},s^{\prime}\}$ is a profitable deviation; (ii) if $s^{\prime}=\overline{p}$, a signal supported on $\overline{p}$ only is a profitable deviation; (iii) $s^{\prime}<\overline{p}$, a signal supported on $\{s^{\prime},s^{\prime\prime}\}$ is a profitable deviation.
B.1 Proof of Claim 4
@@
We first prove the ‘affine’ part. Let $\underline{s}\equiv\min\operatorname*{supp}\pi_{S}$ and $\overline{s}\equiv\max\operatorname*{supp}\pi_{S}$. Since the sender acquires information, $\underline{s}<\overline{p}<\overline{s}$. Let constants $a$ and $b$ be defined by $w_{\beta}(\underline{s})=a+b\underline{s}$ and $w_{\beta}(\overline{s})=a+b\overline{s}$. Suppose that there is some $s\in\operatorname*{supp}\pi_{S}$ with $w_{\beta}(s)<a+bs$. A signal $\pi^{\prime}$ with $\pi_{S}^{\prime}(s)=0$ and strictly higher expected payoff then exists (obtained by appropriately shifting probability mass from $s$ to $\{\underline{s},\overline{s}\}$). Suppose next that there is some $s\in\operatorname*{supp}\pi_{S}$ with $w_{\beta}(s)>a+bs$. A signal $\pi^{\prime}$ with $\pi_{S}^{\prime}(s)>\pi_{S}(s)$ and strictly higher expected payoff then exists (obtained by appropriately shifting probability mass from $\{\underline{s},\overline{s}\}$ to $s$).
@@
We now turn to the ‘strictly increasing’ part. Suppose first that $w_{\beta}$ is constant on $\operatorname*{supp}\pi_{S}$. It must then equal $v(\overline{p})$, since if it were strictly larger (smaller) the sender would be inducing a posterior belief strictly above (below) $\overline{p}$ with probability $1$ in the receiver (since $v$ is non-decreasing). This would contradict that the receiver uses Bayes’ rule to update her belief on path. If $w_{\beta}$ is constant on $\operatorname*{supp}\pi_{S}$, then $w_{\beta}(\underline{s})=v(\overline{p})$. Since the sender can prove news better than the prior, there exists some $m_{h}$ such that $v(\min M^{-1}(m_{h}))>v(\overline{p})$ and, since $v$ is non-decreasing, we also have that $s_{h}\equiv\min M^{-1}(m_{h})>\overline{p}$. So a signal supported on $\{\underline{s},s_{h}\}$ exists and yields an expected payoff strictly above $v(\overline{p})$, constituting therefore a profitable deviation.
Suppose next that $w_{\beta}(s^{\prime})<w_{\beta}(s)$ for some $s^{\prime}>s$, for $s,s^{\prime}\in\operatorname*{supp}\pi_{S}$. Consider any $m^{\prime}\in\operatorname*{supp}\mu(\cdot|\pi,s^{\prime})$ and any $m\in\operatorname*{supp}\mu(\cdot|\pi,s)$. Since $v$ is non-decreasing, it must be that $\beta(m^{\prime})<\beta(m)$. Since the receiver uses Bayes’ rule following on-path messages, it must also be that $m^{\prime}\in\operatorname*{supp}\mu(\cdot|\pi,\tilde{s}^{\prime})$ for some $\tilde{s}^{\prime}<s^{\prime}$ in $\operatorname*{supp}\pi_{S}$, or that $m\in\operatorname*{supp}\mu(\cdot|\pi,\tilde{s})$ for some $\tilde{s}>s$ in $\operatorname*{supp}\pi_{S}$, or both. In all cases, there exists a pair of on-path sender types with the same value of $w_{\beta}$. Since we established that $w_{\beta}$ is affine on $\operatorname*{supp}\pi_{S}$, it must be constant on $\operatorname*{supp}\pi_{S}$, contradicting that $w_{\beta}(s^{\prime})<w_{\beta}(s)$ for some $s^{\prime}>s$, for $s,s^{\prime}\in\operatorname*{supp}\pi_{S}$.

### v6

@@ A.2.2 Proof of Lemma 1
We make use of the following intermediate claim, proved after the main argument is complete.
@@
[…]
Suppose next that $s_{+}$ is not lowest-consistent with any message. Then there exists a $\tilde{s}<s_{+}$ such that $w_{\beta}(\tilde{s})\geq w_{\beta}(s_{+})$. Observe next that the definition of $s_{-}$ and $s_{+}$ and Claim 3 imply that $(\operatorname*{cav}w_{\beta})(s_{+})=w_{\beta}(s_{+})>w_{\beta}(s_{-})=(\operatorname*{cav}w_{\beta})(s_{-})$. Since $w_{\beta}(\tilde{s})\geq w_{\beta}(s_{+})$, we have that
@@
There are three possibilities. If $\tilde{s}<s_{-}$, then $\operatorname*{cav}w_{\beta}$ is not concave, a contradiction. If $\tilde{s}=s_{-}$ then $(\operatorname*{cav}w_{\beta})(s_{-})>(\operatorname*{cav}w_{\beta})(s_{-})$, also a contradiction. If $\tilde{s}\in(s_{-},s_{+})$ then $\operatorname*{cav}w_{\beta}$ is not affine over $s_{-},s_{+}]$, contradicting Lemma 4. It follows that both $s_{-}$ and $s_{+}$ are lowest-consistent with some message, completing the proof of the claim. ∎
@@
Suppose otherwise that $w_{\beta}(s_{-})\geq w_{\beta}(s_{+})$. Since the sender can prove news better than the prior, there exists a message $m_{h}$ such that $v(\min M^{-1}(m_{h}))>v(\overline{p})$. Note that $s_{h}\equiv\min M^{-1}(m_{h})>\overline{p}$, since $v$ is non-decreasing. It follows from the definition of $w_{\beta}$ that $w_{\beta}(s_{h})>v(\overline{p})$.
Observe next that $w_{\beta}(s_{-})\leq v(\overline{p})$, since for any $m\in M(s_{-})$, $\beta(m)\leq s_{-}$ (recall that $\beta(m)=\min M^{-1}(m)$, by construction) and therefore $v(\beta(m))\leq v(s_{-})\leq v(\overline{p})$. This, in turn, implies that $w_{\beta}(s_{h})>w_{\beta}(s_{-})$. Hence, since $w_{\beta}(s_{-})\geq w_{\beta}(s_{+})$ by hypothesis, we have that $w_{\beta}(s_{h})>w_{\beta}(s_{+})$ also holds. Summarizing, we have established that $w_{\beta}(s_{h})>w_{\beta}(s_{-})\geq w_{\beta}(s_{+})$.
Recall that, by definition, $(\operatorname*{cav}w_{\beta})(s_{-})=w_{\beta}(s_{-})$ and $(\operatorname*{cav}w_{\beta})(s_{+})=w_{\beta}(s_{+})$. There are three possibilities to consider. If $s_{h}<s_{+}$, $\operatorname*{cav}w_{\beta}$ is not affine over $(s_{-},s_{+})$, contradicting Lemma 4. If $s_{h}=s_{+}$ then $w_{\beta}(s_{+})>w_{\beta}(s_{+})$, a contradiction. If $s_{h}>s_{+}$ then, since
$(\operatorname*{cav}w_{\beta})(s_{h})>(\operatorname*{cav}w_{\beta})(s_{-})\geq(\operatorname*{cav}w_{\beta})(s_{+}),$
$\operatorname*{cav}w_{\beta}$ is not concave, a contradiction. It follows that $w_{\beta}(s_{-})<w_{\beta}(s_{+})$. ∎
This completes the proof of Lemma 6. ∎
@@ Appendix B Proof of Theorem 1
Let $(\pi,\mu,\beta)$ denote an equilibrium which is not unpersuasive and suppose otherwise that there is an $s^{\prime\prime}\in\operatorname*{supp}\pi_{S}$ and a message $m^{\prime\prime}\in\mu(\cdot|\pi,s^{\prime\prime})$ such that, for some $s^{\prime}<s^{\prime\prime}$, $m^{\prime\prime}\in M(s^{\prime})$.
Since $(\pi,\mu,\beta)$ is not unpersuasive, some on-path message $m$ must lead to a payoff $v(\beta(m))\neq v(\overline{p})$ for the sender, meaning that $\beta(m)\neq\overline{p}$. It follows that the sender must be acquiring some information in equilibrium, so that $\operatorname*{supp}\pi_{S}\neq\{\overline{p}\}$. Let $\underline{s}\equiv\min\operatorname*{supp}\pi_{S}$ and $\overline{s}\equiv\max\operatorname*{supp}\pi_{S}$; since $\operatorname*{supp}\pi_{S}\neq\{\overline{p}\}$, $\sum_{s\in\operatorname*{supp}\pi_{S}}\pi_{S}(s)s=\overline{p}$ implies that $\underline{s}<\overline{p}<\overline{s}$. Define the sender’s equilibrium interim value from signal realization $s\in[0,1]$ as $w_{\beta}(s)\equiv\max_{m\in M(s)}v(\beta(m))$.
The argument relies on the following claim, proved after the main argument is complete.
@@
Suppose first that $s^{\prime\prime}\leq\overline{p}$. $w_{\beta}$ being affine (from Claim 4) together with $\pi$ being a feasible signal, meaning that $\sum_{s\in\operatorname*{supp}\pi_{S}}\pi_{S}(s)s=\overline{p}$, imply that the sender’s equilibrium expected payoff can be written as
@@
Consider now a deviating signal supported on $\{s^{\prime},\overline{s}\}$ only. By choosing this signal, the sender obtains an expected payoff equal to
@@
To show that (8) is strictly larger than (7) it suffices to note that: $w_{\beta}(s^{\prime})\geq w_{\beta}(s^{\prime\prime})$, since $m^{\prime\prime}$ is optimal at $s^{\prime\prime}$ and $m^{\prime\prime}\in M(s^{\prime})$, by hypothesis; $s^{\prime}<s^{\prime\prime}$, by hypothesis; and $w_{\beta}(\overline{s})>w_{\beta}(s^{\prime\prime})$, since $w_{\beta}$ is strictly increasing on $\operatorname*{supp}\pi_{S}$, by Claim 4. This contradicts that $\pi$ is an equilibrium signal.
@@
Suppose next that $s^{\prime\prime}>\overline{p}$. There are three cases to consider and, in each, a deviation analogous to the one in the previous paragraph can be constructed: (a) if $s^{\prime}>\overline{p}$, a signal supported on $\{\underline{s},s^{\prime}\}$ is a profitable deviation; (b) if $s^{\prime}=\overline{p}$, a signal supported on $\overline{p}$ only is a profitable deviation; (c) $s^{\prime}<\overline{p}$, a signal supported on $\{s^{\prime},s^{\prime\prime}\}$ is a profitable deviation. This establishes item (i) in the theorem.
To prove item (ii), consider any on-path message $m$. Item (i) implies that message $m$ is sent by type $\min M^{-1}(m)$ only: $m\in\operatorname*{supp}\mu(\cdot|\pi,\min M^{-1}(m))$ and $m\notin\operatorname*{supp}\mu(\cdot|\pi,s)$ if $s\neq\min M^{-1}(m)$. As the receiver’s beliefs are to satisfy Bayes’ rule on the equilibrium path, it must be that $\beta(m)=\min M^{-1}(m)$.
For item (iii), consider any on-path sender type $s\in\operatorname*{supp}\pi_{S}$ and a message $m\in\operatorname*{supp}\mu(\cdot|\pi,s)$. Item (i) implies that $s=\min M^{-1}(m)$ and item (ii) implies that $\beta(m)=\min M^{-1}(m)$. Therefore, $\beta(m)=s$. ∎
Proof of Claim 4.
We first prove the ‘affine’ part. Recall that $\underline{s}\equiv\min\operatorname*{supp}\pi_{S}$, $\overline{s}\equiv\max\operatorname*{supp}\pi_{S}$ and that, since $\operatorname*{supp}\pi_{S}\neq\{\overline{p}\}$, $\sum_{s\in\operatorname*{supp}\pi_{S}}\pi_{S}(s)s=\overline{p}$ implies that $\underline{s}<\overline{p}<\overline{s}$. Let constants $a$ and $b$ be defined by $w_{\beta}(\underline{s})=a+b\underline{s}$ and $w_{\beta}(\overline{s})=a+b\overline{s}$.
@@
Suppose that there is some $s\in\operatorname*{supp}\pi_{S}$ with $w_{\beta}(s)<a+bs$. By appropriately shifting probability mass from $s$ to $\{\underline{s},\overline{s}\}$, we can construct a signal $\pi^{\prime}$ such that $\pi_{S}^{\prime}(s)<\pi_{S}(s)$ which leads to a strictly higher expected payoff for the sender.
Suppose next that there is some $s\in\operatorname*{supp}\pi_{S}$ with $w_{\beta}(s)>a+bs$. By appropriately shifting probability mass from $\{\underline{s},\overline{s}\}$ to $s$ we can construct a signal $\pi^{\prime}$ with $\pi_{S}^{\prime}(s)>\pi_{S}(s)$ which leads to a strictly higher expected payoff for the sender.
@@
We now turn to the ‘strictly increasing’ part. Suppose first that $w_{\beta}$ is constant on $\operatorname*{supp}\pi_{S}$. It must then equal $v(\overline{p})$, since if it were strictly larger (smaller) the sender would be inducing in the receiver a posterior belief strictly above (below) $\overline{p}$ with probability one (since $v$ is non-decreasing). But if $w_{\beta}$ is constant and equal to $v(\overline{p})$ on $\operatorname*{supp}\pi_{S}$, the equilibrium must be unpersuasive, which does not hold by hypothesis.
[…]
Appendix C Proof of Theorem 2
Start with the first part of the statement: we want to show that if PNBP holds then all equilibria are unraveling. Exploiting Theorem 1, it is enough to show that if PNBP holds and an equilibrium is unpersuasive, then it is also an unraveling equilibrium. Let $(\pi,\mu,\beta)$ denote such an unpersuasive equilibrium. Observe that, since PNBP holds, there exists some $m_{h}$ such that $v(\min M^{-1}(m_{h}))>v(\overline{p})$ and, since $v$ is non-decreasing, we also have that $s_{h}\equiv\min M^{-1}(m_{h})>\overline{p}$.
Suppose first that the sender acquires information, so that $\operatorname*{supp}\pi_{S}\neq\{\overline{p}\}$. Let $\underline{s}\equiv\min\operatorname*{supp}\pi_{S}$. Since the sender acquires information, $\underline{s}<\overline{p}$. Since the equilibrium is unpersuasive, we have that $v(\beta(m))=v(\overline{p})$ for every $m\in\operatorname*{supp}\mu(\cdot|\pi,\underline{s})$. Then the sender can profitably deviate by choosing signal $\pi^{\prime}$ with $\operatorname*{supp}\pi_{S}^{\prime}=\{\underline{s},s_{h}\}$. So, when PNBP holds, there cannot be unpersuasive equilibria in which the sender acquires information.
Consider now the case in which the sender acquires no information in equilibrium, so that $\pi_{S}(\overline{p})=1$. Suppose that there exists some $m^{\prime\prime}\in M(\overline{p})$ such that $\mu(m^{\prime\prime}|\pi,\overline{p})>0$ for which we can find some $s^{\prime}<\overline{p}$ such that $m^{\prime\prime}\in M(s^{\prime})$. That is: $\overline{p}$ is not lowest-consistent with message $m^{\prime\prime}$.
We will show that the sender can profitably deviate by choosing signal $\pi^{\prime}$ with $\operatorname*{supp}\pi_{S}^{\prime}=\{s^{\prime},s_{h}\}$. Following realization $s^{\prime}$, the sender obtains a payoff of at least $v(\overline{p})$ (since $m^{\prime\prime}\in M(s^{\prime})$, by construction) and following realization $s_{h}$ the sender obtains a payoff strictly above $v(\overline{p})$ (since $m_{h}\in M(s_{h})$, $\beta(m_{h})\geq s_{h}$ and therefore $v(\beta(m_{h}))\geq v(s_{h})>v(\overline{p})$). Hence $\pi^{\prime}$ is a strictly profitable deviation, so if $(\pi,\mu,\beta)$ is an equilibrium in which the sender acquires no information, it must be that $\overline{p}$ is lowest-consistent with every $m\in\operatorname*{supp}\mu(\cdot|\pi,\overline{p})$. It follows that $\beta(m)=\overline{p}=\min M^{-1}(m)$ for every $m\in\operatorname*{supp}\mu(\cdot|\pi,\overline{p})$, so that the receiver is maximally skeptical following every on-path message and that the sender reveals all acquired information. The equilibrium is therefore unraveling.
Now to the second part of the statement. Such an unpersuasive equilibrium is constructed in Lemma 5, Appendix A. Note that Lemma 5 does not rely on the continuity requirement stated in Lemma 1, which is not assumed for Theorem 2.
To show that it is sender-preferred, we will show that there is no equilibrium in which the sender obtains an expected payoff strictly larger than $v(\overline{p})$. Suppose that $(\pi,\mu,\beta)$ is such an equilibrium. It cannot be unpersuasive by definition, so Theorem 1 implies that it must be unraveling. Therefore, Corollary 1 implies that the sender’s expected payoff is equal to $(\operatorname*{cav}\underline{v}_{M})(\overline{p})$, where $\underline{v}_{M}$ is defined in (3). However, if $(\operatorname*{cav}\underline{v}_{M})(\overline{p})>v(\overline{p})$ there must exist some $s\in[0,1]$ such that $\underline{v}_{M}(s)>v(\overline{p})$, implying in turn that there exists an $m_{h}\in\mathcal{M}$ such that $v(\min M^{-1}(m_{h}))>v(\overline{p})$, i.e., that the sender can prove news better than the prior, a contradiction. ∎
Appendix D Proof of Theorem 3
We prove first Lemma 2 and then turn to the proof of Theorem 3. Recall that in Section 4 we are maintaining the condition stated in Lemma 1, which ensures equilibrium existence.
D.1 Proof of Lemma 2
Start with the $\Rightarrow$ direction. We use a replication argument which exploits Theorem 1. Consider first the triple $(v,\overline{p},M^{\prime})$ and any unraveling equilibrium signal $\pi^{\prime}$, which exists by hypothesis and leads to payoff $\underline{V}_{M^{\prime}}^{*}(v,\overline{p})$. This is because, if PNBP holds, every equilibrium is unraveling (by Theorem 2) and the sender’s expected payoff is unique (by Corollary 1). If PNBP does not hold, the unraveling equilibrium expected payoff does not exceed the unpersuasive equilibrium payoff, from Theorem 2.
We will show that, in any equilibrium with parameters $(v,\overline{p},M^{\prime\prime})$, the sender’s expected payoff from $\pi^{\prime}$ is at least as large as $\underline{V}_{M^{\prime}}^{*}(v,\overline{p})$. This, in turn, means that the expected payoff in any equilibrium with parameters $(v,\overline{p},M^{\prime\prime})$ must also be at least as large as $\underline{V}_{M^{\prime}}^{*}(v,\overline{p})$, completing the argument.
Observe that in any unraveling equilibrium at parameters $(v,\overline{p},M^{\prime})$, Theorem 1 implies that for any equilibrium signal $\pi^{\prime}$ at any $s\in\operatorname*{supp}\pi_{S}^{\prime}$ the sender obtains a continuation payoff of $v(s)$. Since $s$ is lowest-consistent with some message under $M^{\prime}$ and $L_{M^{\prime\prime}}\supseteq L_{M^{\prime}}$, $s$ is also lowest-consistent with some message under $M^{\prime\prime}$. This in turn implies that, for any equilibrium (unraveling or not) $(\pi^{\prime\prime},\mu^{\prime\prime},\beta^{\prime\prime})$ when parameters are $(v,\overline{p},M^{\prime\prime})$, $\max_{m\in M^{\prime\prime}(s)}v(\beta^{\prime\prime}(m))\geq v(s)$ for every $s\in\operatorname*{supp}\pi_{S}^{\prime}$. Hence the sender’s expected payoff from $\pi^{\prime}$ with parameters $(v,\overline{p},M^{\prime\prime})$ is at least as large as $\underline{V}_{M^{\prime}}^{*}(v,\overline{p})$, as required.
Now to the $\Leftarrow$ direction. We will prove the contrapositive, i.e., that $M^{\prime\prime}\not\succeq^{\mathrm{lc}}M^{\prime}$ implies $\underline{V}_{M^{\prime\prime}}^{*}(v,\overline{p})<\underline{V}_{M^{\prime}}^{*}(v,\overline{p})$ for some admissible $v$ and $\overline{p}$ such that at $(v,\overline{p},M^{\prime})$ an unraveling equilibrium exists. Observe first that $M^{\prime\prime}\not\succeq^{\mathrm{lc}}M^{\prime}$ implies that $L_{M^{\prime}}\neq\{0\}$, as $0\in L_{M}$ for any $M$. Since $L_{M^{\prime}}\neq\{0\}$, there exists some $s^{*}\in L_{M^{\prime}}$, $s^{*}>0$ such that $s^{*}\notin L_{M^{\prime\prime}}$.
Take $v(p)=1_{\{p\geq s^{*}\}}$ and $\overline{p}=s^{*}/2$. Let $\underline{v}_{M^{\prime}}$ and $\underline{v}_{M^{\prime\prime}}$ denote the value of posteriors with maximal receiver skepticism as defined in equation (3) when the verifiability structures are $M^{\prime}$ and $M^{\prime\prime}$, respectively. With parameters $(v,\overline{p},M^{\prime})$ PNBP holds, so all equilibria are unraveling (Theorem 2) and lead the sender to obtain an expected payoff equal to her full-commitment payoff of $(\operatorname*{cav}v)(\overline{p})=1/2$ (this follows from observing that $(\operatorname*{cav}v)(\overline{p})=(\operatorname*{cav}\underline{v}_{M^{\prime}})(\overline{p})$ and applying Corollary 1).
This expected payoff is not attainable in any equilibrium under $M^{\prime\prime}$. If the sender cannot prove news better than the prior under $M^{\prime\prime}$, this follows from Theorem 2, as the largest expected payoff for the sender in equilibrium is $v(\overline{p})=0<1/2$ in this case. If the sender can prove news better than the prior under $M^{\prime\prime}$, this follows because all equilibria must be unraveling (from Theorem 2), so $s^{*}\notin L_{M^{\prime\prime}}$ implies that $s^{*}$ cannot be on-path. However, in an unraveling equilibrium, $s^{*}$ must be on-path if the sender is to obtain an expected payoff of $(\operatorname*{cav}v)(\overline{p})$. ∎
D.2 Proof of Theorem 3
For the $\Rightarrow$ direction, observe that if at $(v,\overline{p},M^{\prime})$ the sender can prove news better than the prior, then an unraveling equilibrium exists (Theorem 2) and the equilibrium expected payoff is unique (Corollary 1), i.e., $\underline{V}_{M^{\prime}}^{*}(v,\overline{p})=\overline{V}_{M^{\prime}}^{*}(v,\overline{p})=(\operatorname*{cav}\underline{v}_{M^{\prime}})(\overline{p})$. Lemma 2 therefore implies that $\underline{V}_{M^{\prime\prime}}^{*}(v,\overline{p})\geq\overline{V}_{M^{\prime}}^{*}(v,\overline{p})$. As $M^{\prime\prime}\succeq^{\mathrm{lc}}M^{\prime}$, the sender can also prove news better than the prior at $(v,\overline{p},M^{\prime\prime})$, so that $\underline{V}_{M^{\prime\prime}}^{*}(v,\overline{p})=\overline{V}_{M^{\prime\prime}}^{*}(v,\overline{p})=(\operatorname*{cav}\underline{v}_{M^{\prime\prime}})(\overline{p})$ (from Theorem 2 and Corollary 1), completing this part of the argument.
Now to the $\Leftarrow$ direction. We will prove the contrapositive, i.e., that $M^{\prime\prime}\not\succeq^{\mathrm{lc}}M^{\prime}$ implies $\overline{V}_{M^{\prime\prime}}^{*}(v,\overline{p})<\overline{V}_{M^{\prime}}^{*}(v,\overline{p})$ for some admissible $v$ and $\overline{p}$ such that at $(v,\overline{p},M^{\prime})$ PNBP holds. Observe that the construction used in the proof of Lemma 2 in fact considers a triple $(v,\overline{p},M^{\prime})$ such that the sender can prove news better than the prior and shows that the unique equilibrium expected payoff at $(v,\overline{p},M^{\prime})$ is not attainable in any equilibrium under $M^{\prime\prime}$. ∎